OpenStudy (anonymous):

Please show me the work-an equation of the line that is perpendicular to the given line -10x + 8y = 3 and passes through the given point (15, 12).

OpenStudy (anonymous):

8y=10x+3

OpenStudy (anonymous):

y=10/8x+3/8

OpenStudy (anonymous):

now take the slope and plug it into the equation \[y-y_o=m(x-x_o)\]

OpenStudy (anonymous):

\[y-12=\frac{10}{8}(x-15)\]

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

you could probably simplify it a bit more if you wanted to

OpenStudy (anonymous):

how would I goabout that?

OpenStudy (anonymous):

How do I graph it

OpenStudy (anonymous):

y=1.25x-6.75

OpenStudy (anonymous):

I would then pick a number for x try zero and one. at point (0,-6.75) and at (1,-5.5) you can then plot those two points on a graph and connect the dots